OpenStudy (anonymous):

A jet boat is used to take canoers 200 miles upriver to their starting point. As the driver of the jet boat, Jeremy knows that it will take him 5 hours to get the canoers upstream, but only 4 hours to get the boat back downstream after dropping the canoers off, as long as he keeps the boat moving at a constant speed. How fast would the boat be moving in still water? A. 5 mph B. 40 mph C. 45 mph D. 50 mph

OpenStudy (ranga):

Effective speed when going upstream is: 200 miles / 5 hours = 40 mph. Effective speed when going downstream is: 200 miles / 4 hours = 50 mph. Let u be the speed of the boat and r be the speed of the river. When going upstream, the effective speed of the boat = u - r = 40 When going downstream, the effective speed of the boat = u + r = 50 Solve for u from these two equations.